an object travels in a straight line at 55 mph for three minutes and 23 seconds how many meters did it travel?
Answer by mananth(12270)

3 minutes 23 seconds =
180+23= 203 seconds
1 hour = 3600 seconds
203 seconds = 203/3600 hours
D= r*t
D= 55* (203/3600)
D=3.10 miles
1 mile = 1600 meters
3.10 miles*1600
=4960 meters